Dec 3, 2015 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Update BCAL clusterizer to improve agreement between simulated and generated MC (Mark)
  2. Modify covariance matrix for FCAL and BCAL to correspond to the current resolution of the detectors (Matt).
  3. Change F250 simulation code to match firmware and add flexibility for new algorithms (Mike)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Preparations
    1. FCAL Update
    2. BCAL Update
  4. Reconstruction and Simulation
  5. Calibration
    1. Calibration Challenge
    2. Layer Efficiency in Mod 7/8 Runs [3](Ahmed)
  6. Any other business


Attendees: Matt, David, Tegan (JLab), Mike, Will (CMU), Christina, George (Athens), Sean (NWU), Andrei, Noemi, Ahmed, Zisis (Regina) "

  1. Announcements: shifts cancelled until 4pm on Thu. Accelerator is seeking a solution on magnet problem to give us full energy (11 GeV in Hall A).
  2. Action Items: Mark not here. Matt: nothing to report. Mike (emulation): procedure for future data seems straight forward, but to implement for old data it is tricky. Not complicated, but things need to be parsed out, from config files to actual analysis. Need two sets of calibrations for raw data and processed. NSA and NSB need to be passed around in a new way; they should be in RCDB database and obtained by the emulation software from there ((we had NSA+NSB but not the values separately). Begin-of-run record will be available for spring data. Go after new data solution first and probably not worry about old data. Bring up such discussions in a f250-type meeting so offset folks can participate (Elton to organize?).
  3. Run Preparations
    1. FCAL Update: hot checkout completed. Lookout for noisy channels in the trigger (talk to Sasha) - problem was traced to coax cable sheath broken by firm (serial no range is known for this bases). Heavens is implementing alarm channels.
    2. BCAL Update:
      1. Hot checkout done, detector running.
      2. Noemi: progressing, something to show next week on time-walk.
      3. Tegan: Putting in another hits factory to implement sampling fractions, since those are based on energy and angle. Plan is to construct IU showers, go back and correct for SF.
      4. George: nothing to report; things on hold. University studies.
  4. Reconstruction and Simulation
    • FCAL Hadron reconstruction Log Entry 3360676 (Elton): Wait until Elton comes back.
    • FCAL MC nonlinear corrections Log Entry 3361053 (Elton): Matt looked at these. The function in Mihajlo's technical not, it looks coded correctly, but factor ranges from 0.5-2 over a few GeV. Nothing has changed since then. There is an issue, but there is no understanding yet. Equation can not be solved in its current form, but rather needs iteration.
    • BCAL low energy pi0s [4] (Will M) Low energy pi0 discussed in physics meeting. Inclusive pi0s: Peak starts appearing between 50-60 MeV, which agrees with our (ancient) expectations. Exclusive pi0s (lower stats): it appears at 75 MeV but low in mass. Might have non-linearity at low energies explaining to low pi0 mass value.
  5. Calibration
    1. Calibration Challenge: Still in testing phase, first results probably next week. Talking to Adesh to integrate his code. BCAL things are stable. On simulation side, things are setup to run large scale bggen simulations. Paul's p vs beta for pions in BCAL does not quite get to 1 at high momenta. This is not quite understood yet.
    2. Layer Efficiency in Mod 7/8 Runs [5](Ahmed): Follow up on Mod 7/8 efficiencies. Difference appears to be caused by the emulations. Difference between Mod 7/8 is not difference based on the three variable studied here: no of hits per run, point energies, point time, integral samples. David: change config parameters to match the NSA+NSAB=55 to 60 samples for Mod 8. (Nothing can be done about threshold obviously.) Ahmed will investigate.
  6. Any other business: nothing.